# Hybrid satellite-fiber quantum network

### v1

@@ Abstract
Quantum networks hold promise for key distribution, private and distributed computing, and quantum sensing, among other applications. The scale of such networks for ground users is currently limited by one’s ability to distribute entanglement between distant locations. This can in principle be carried out by transmitting entangled photons through optical fibers or satellites. The former is limited by fiber optic attenuation while the latter is limited by atmospheric extinction, diffraction. Here, we propose a hybrid network and protocol that outperform both ground- and satellite-based designs and lead to high-fidelity entanglement at a continental or even global scale.
@@
The ongoing development of quantum technologies calls for advances in quantum communication. The basis of quantum communication is entanglement, which, once established, can allow the transmission of quantum information through arbitrary distances [1]. Multiple small-scale quantum networks have been successfully implemented [2, 3, 4, 5, 6]. The prospect of developing large-scale quantum networks is motivated by applications ranging from quantum key distribution [7, 8] and distributed quantum computing [9] to quantum navigation and sensing [10, 11]. However, scalability remains a challenge due to difficulties associated with the generation of entanglement between distant parties, and thus new approaches are needed to lay the ground for the future development of a quantum internet [12, 13].
@@
The most direct approach for distributing entanglement in a quantum network is by using optical fibers to transmit entangled photons [14, 15]. Due to the exponential loss of photons with distance, establishing long-distance entanglement requires intermediate quantum repeaters [16, 17]. Such devices, which are currently under development, are able to generate entangled states, transmit/receive entangled photons, perform quantum operations, and/or store qubits. Within this approach, entanglement is first directly distributed within short distances. Long-distance entanglement is then established through entanglement swapping performed at intermediate quantum repeaters. However, existing quantum repeater technology currently limits entanglement distribution over long distances [18].
@@
An alternative approach is to use satellites as sources of entangled photon pairs that are sent to two distant detection stations on the ground [19]. Entanglement distribution is successful if both ground stations detect photons of the same pair [20]. The challenge with this approach is that photons are subject to diffraction loss, where the survival probability scales with the inverse square of the distance, and absorption and scattering in Earth’s atmosphere. In 2016, China launched Micius, the first quantum communication low Earth orbit (LEO) satellite [21], which was equipped with an entangled photon pair source and operated in an orbit 500 km from the ground. Entanglement distribution was established between two ground stations 1200 km apart from each other [20]. Moreover, this LEO satellite was successfully used to allow quantum key distribution (QKD) between China and Europe that are 7600 km apart on Earth as a trusted relay [22]. Subsequent work has also developed scheduling policies for entanglement distribution using groups of LEO satellites [23].
@@
While using LEO satellites to distribute entanglement results in low photon loss, certain challenges with this technology currently impact its scalability [21]. First, LEO satellites travel at high speeds relative to the ground, leaving only short time windows for interaction with the corresponding stations, and thus require high precision acquisition, pointing, and tracking systems. Second, the short distances between LEO satellites and the ground limit the area that can be covered by a single satellite, making it impossible for a satellite to establish entanglement between stations far apart. This second shortcoming can be addressed by allowing the satellite to temporarily store qubits as a trusted node, but this could compromise security [24]. An alternative to mitigate both shortcomings is to use geostationary satellites, positioned at 36 000 km above the equator in a fixed position relative to the ground [25]. A single geostationary satellite can cover more than one-third of the Earth’s surface, but the loss caused by diffraction is $10^{4}$ times higher than for Micius, which limits the entanglement distribution rate.
@@
In this article we investigate a hybrid quantum network consisting of a combination of optical fibers and medium Earth orbit (MEO) satellites. The goal is to benefit from the short-distance advantage of optical fibers and the long-distance advantage of satellites. Our focus on MEO satellites—which is chosen to be 10 000 km above ground—seeks to strike a compromise between low photon loss and high spatio-temporal coverage. Using the contiguous United States as a model system, we show that the hybrid protocol provides better large-scale quantum communication performance than both optical-fiber-based and satellite-based protocols.
@@
II II. Optical fiber transmission
Because the number of photons successfully transmitted through an optical fiber decreases exponentially with distance, various quantum repeater designs have been proposed to compensate for the loss [17]. In particular, a promising trapped-ion quantum repeater has been recently demonstrated in an entanglement distribution experiment between points separated by 50 km of optical fibers [26]. A trapped-ion repeater includes at least two trapped ions and the ability to emit photons entangled with each ion. The trapped-ions in the repeaters serve as quantum memories storing qubits. Entanglement swapping between the ions can then be implemented through deterministic Bell state measurement (DBSM).
The ground portion of our hybrid quantum network will consist of a natural extension of this scheme in which long-distance entanglement distribution is achieved using multiple quantum repeaters [Fig. 1(a)]. In this extended scheme, a photon repeater is placed between every pair of trapped-ion repeaters (including the end nodes) [26]. The photon repeaters can detect photons and perform quantum entanglement swapping using probabilistic Bell state analysis. Alice and Bob (the end nodes) are assumed to have one trapped-ion and, as trapped-ion repeaters, the ability to emit ion-entangled photons. Thus, the system consists of two end nodes, $n_{\text{p}}$ photon repeaters, $n_{\text{i}}=n_{\text{p}}-1$ trapped-ion repeaters for $n_{\text{p}}\geq 1$ ($n_{\text{i}}=0$ otherwise). In this system, ion-photon entangled states $\rho_{\text{ip}}$ are converted into ion-ion entangled states $\rho_{\text{ii}}$a, and then to an end-to-end entangled state $\rho_{\text{AB}}$ shared by Alice and Bob.
Figure 2: Entanglement distribution rate and fidelity through optical fibers. (a) Distribution rate without repeaters (black) and for an optimal number of evenly-placed photon (and thus intermediate trapped-ion) repeaters (blue) as a function of the optical fiber length $L$. The optimal number of photon repeaters is marked on the top axis. Inset: magnification of both curves for small $L$. (b) Resulting end-to-end fidelity as the number of evenly-placed repeaters is varied. (c) Corresponding entanglement distribution rate for two fixed values of $L$ as the number of repeaters is varied.
@@
where $L$ is the fiber length between the end nodes, and the other symbols are defined as follows. The parenthesis accounts for the time $\tau$ needed to emit an entangled photon and the time to transmit the photon to an adjacent photon repeater, where $c=2\times 10^{8}$ m/s is the speed of light in the optical fiber; note that a factor 1/2 is canceled because the trapped-ion nodes wait to hear back from photon repeaters whether the photon was detected or not before attempting to send a new photon. The term $\frac{1}{2}P^{2}\eta^{2}$ is the probability that both photons concurrently directed to a photon repeater are successfully received, detected, and projected into a Bell state; here, $\eta(L)=10^{-\gamma L/2n_{\text{p}}}$ is the fiber transmission rate for an attenuation rate $\gamma$, $P$ is the photon detector efficiency, and 1/2 is the probability of a successful photon Bell state measurement. The waiting time to establish a successful entanglement between one pair of adjacent trapped-ion nodes follows the geometric distribution; thus, the additional factor $(3/2)^{\nu}$ is a good approximation for the waiting time before all pairs in the path are ready to be entangled, where $\nu=\log_{2}{n_{\text{p}}}$. The time for local quantum operations associated with entanglement swapping is assumed to be negligible and with a success rate of 1/2 for photon repeaters (as mentioned) and 1 for trapped-ion repeaters. The average entanglement distribution rate is then given by $R_{\text{f}}=1/T_{\text{f}}$. As a reference, we note that the average time to establish an entanglement between Alice to Bob without any intermediate repeater is
@@
where $\eta^{\prime}(L)=10^{-\gamma L}$, and the average entanglement distribution rate thus given by $R_{\text{f}}^{\prime}=1/T_{\text{f}}^{\prime}$.
Figure 2(a) shows the resulting distribution rate $R_{\text{f}}$ for the optimal number of repeaters $n_{\text{p}}^{*}$ as the end-to-end fiber length $L$ is varied, obtained by minimizing $T_{\text{f}}$ as a function of $n_{\text{p}}$ in Eq. (1). We also show the distribution rate $R_{\text{f}}^{\prime}$, which is higher for distances smaller than $61.7$ km (inset). Note that at this distance, the optimal number of photon repeaters transitions from 0 to 7 (rather than to 1); this gap comes from the competition (as $n_{\text{p}}$ is increased) between the increase in the entanglement distribution rate with neighbors and the decrease in the probability of successful Bell state measurements for all photon repeaters in the path. Following [26], our simulations assume that photons are emitted entangled with 40Ca+ ions and converted to 1550 nm, which corresponds to the lowest attenuation in the optical fiber, $\gamma=0.0173$ km-1. The other parameters are $P=0.21$ and $\tau=175$ $\mu$s, where the efficiency of photon conversion is incorporated in $P$. At large distances, repeaters allow the decay of the entanglement distribution rate to be substantially slower than the exponential one observed without repeaters, but the optimal number of repeaters increases linearly with distance. Note that for a fixed fiber length, the entanglement distribution rate increases monotonically as $n_{\text{p}}$ increases before reaching the optimum at $n_{\text{p}}^{*}$ [Fig. 2(c)].
A competing factor to be accounted for is the end-to-end fidelity as the number of repeaters increases. We follow the model discussed in [28] to calculate the fidelity as a function of $n_{\text{p}}$. We assume that the initially generated ion-photon entanglement is a mixed state of the form $\rho_{\text{ip}}=F_{0}|\Psi^{-}\rangle\langle\Psi^{-}|+(1-F_{0})|\Psi^{+}\rangle\langle\Psi^{+}|$, where $|\Psi^{\pm}\rangle$ are Bell states and $F_{0}$ is the pairwise fidelity [Fig. 1(a)]. A successful Bell state measurement on a photon pair, which occurs with a probability $1/2$, leads to a new mixed entangled state $\rho_{\text{ii}}=F_{\text{ii}}|\Psi^{-}\rangle\langle\Psi^{-}|+(1-F_{\text{ii}})|\Psi^{+}\rangle\langle\Psi^{+}|$ shared by two neighboring trapped-ion nodes. The ion-ion entanglement fidelity is $F_{\text{ii}}=[1+V(1-2F_{0})^{2}]/2$, where $V=2F_{\text{swap}}^{\text{p}}-1$ is the photon interference visibility. Then each intermediate trapped-ion repeater node can perform a DBSM on the two trapped-ion qubits they maintain. Given $\rho_{\text{i}_{1}\text{i}_{2}}$ and $\rho_{\text{i}_{3}\text{i}_{4}}$, an ideal DBSM will give $\rho_{\text{i}_{1}\text{i}_{4}}$ by projecting trapped-ions 2 and 3 on to a Bell state and tracing them out. If we account for the depolarizing noise caused by the DBSM, the resulting quantum state will be $\tilde{\rho}_{\text{i}_{1}\text{i}_{4}}(\rho_{\text{i}_{1}\text{i}_{4}},F^{\text{i}}_{\text{swap}})$, where $F_{\text{swap}}^{\text{i}}$ is the trapped-ion entanglement swapping fidelity and
@@
Here, $S_{j}\in[I\otimes\sigma_{x},I\otimes\sigma_{y},I\otimes\sigma_{z}]$, where $\sigma_{j}$ are the Pauli matrices and $I$ is the identity matrix. The fidelity of the final end-to-end entangled state is obtained after completing all the entanglement swapping operations on the intermediate trapped-ion nodes. Through the entire process, the fidelity of the established entanglement can decay due to the decoherence of qubits stored in the quantum memories. For an arbitrary initial state $\rho_{0}$, this can be modeled by considering dephasing noise in the form
@@
where $q(t)=(1-e^{-t^{2}/\tau_{q}^{2}})/2$ and $\tau_{q}$ is the decoherence time. A recent study shows that the decoherence time of trapped-ion entanglement can be longer than 10 s [29], which is much longer than the relevant time needed to establish one end-to-end entanglement discussed in this study. Therefore, we assume the impact of decoherence to be negligible in our calculations, leading to no fidelity loss. An analysis of the impact of non-negligible decoherence is included in the Supplemental Material [39].
The end-to-end entanglement fidelity $F_{\text{f}}$ decreases as the number of quantum repeaters increases. Entanglement is completely lost for $n_{\text{p}}>26$ when the three parameters are $F_{0}$ = $F_{\text{swap}}^{\text{p}}$ = $F_{\text{swap}}^{\text{i}}$ = 0.99 (as assumed throughout unless otherwise suggested) [Fig. 2(b), blue] [30]. Therefore, entanglement distillation is necessary to achieve high end-to-end fidelity, which we implement using the recurrence method for mixed-state distillation. Given two entangled mixed states with fidelities $F_{1}$ and $F_{2}$ shared by the same pair of nodes, a successful purification yields a single entanglement with a higher fidelity $F_{\text{p}}=(10F_{1}F_{2}-F1-F2+1)/(8F_{1}F_{2}-2F_{1}-2F_{2}+5)$. The probability of completing such a 2$\rightarrow$1 purification successfully is $(8F_{1}F_{2}-2F_{1}-2F_{2}+5)/9$. With multiple copies of entanglement shared by two neighboring nodes, the 2$\rightarrow$1 purification can be iteratively repeated until only one entangled state remains for the given pair of nodes. The distillation process will allow us to reduce the optimal number of repeaters used in Fig. 2, which was generated without considering purification.
Figure 3: Model optical-fiber network of the contiguous U.S. (a) Network structure, where each node represents a census tract and each link represents an optical fiber. Links that require intermediate repeaters, namely those longer than 61.7 km, are marked green. The orange path illustrates an entanglement distribution route between end nodes Alice and Bob. (b-d) Histograms of node degrees (b), link lengths (c), and optical-fiber path lengths between 10 0001000010\,00010 000 randomly sampled end node pairs (d) in the network.
@@
We consider a MEO satellite orbiting 10 000 km above ground. The satellite is equipped with an entangled-photon source that can generate $N_{\text{s}}$ entangled photon pairs per second. The photon pairs are split into two collimated beams and are then sent toward two ground stations as shown in Fig. 1(b). An entanglement is successfully established between two ground stations if both stations detect photons from the same entangled pair. The ground stations will then store the quantum states of the received photons in single-photon quantum memories [which also use trapped ions, but are different from the trapped-ion memories that operate by emitting photons in Fig 1(a)] [31]. Free space diffraction and atmospheric extinction (absorption and scattering) are accounted for as the main factors causing losses during transmission [32]. Other losses come from limitations in the transmitting and receiving telescopes.
Figure 4: Entanglement distribution efficiency through the exclusive use of optical fibers or MEO satellites for 10 000 randomly sampled Alice and Bob pairs in the network shown in Fig. 3(a). (a) Distribution rate through optical fibers (blue) and a satellite (orange). For a fair comparison, the fidelity for the satellite and target fidelity for the ground network are both assumed to be $0.87$. (b) Fidelity of the end-to-end entanglement for the time needed to establish one entanglement via satellite, color-coded as in (a). The points $D^{\star}_{\text{R}}$ and $D^{\star}_{\text{F}}$ mark the distances where the relative efficiency inverts. Repeaters are placed to achieve optimal distribution rate in (a) and optimal fidelity in (b). (c) Scattered plot (purple) and histogram (grey) of the distribution rate in (a) and fidelity in (b).
@@
$w_{\text{d}}(z)=w_{0}\sqrt{(1+z^{2}/z_{R}^{2})},$ (5)
where $z_{R}=\pi w_{0}^{2}\lambda^{-1}$ is the Rayleigh range, $\lambda$ is the wavelength, and $w_{0}$ is the initial beam width. The slant range $z$ can be easily calculated given the longitudes and latitudes of the ground station and the subsatellite point (the projection point of the satellite on the Earth’s surface) by converting them to Cartesian coordinates. Given a telescope receiver of a radius $a_{\text{r}}$, the fraction of the initial beam that can be detected is
@@
for the far field $z\gg z_{R}$ relevant for MEO satellites. Absorption and Rayleigh/Mie scattering occur when the beam travels through the atmosphere (mainly the $30$ km layer above ground). The fraction of the beam surviving absorption and scattering is
@@
where $\eta(z,\theta)$ is the photon transmission rate between the satellite and one ground station. This rate accounts for the efficiency $\eta_{\text{tran}}$ of the transmitting telescope and the optics efficiency $\eta_{\text{rec}}$, detection efficiency $\eta_{\text{dec}}$, and spectral filter efficiency $\eta_{\text{filt}}$ of the receiving telescope at the ground station [33, 34]. The photons also need to be converted to a wavelength and spatio-temporal mode compatible with the quantum memories at the ground stations. Since the conversion efficiency can vary depending on the type of memory used, we assume a perfect conversion for simplicity; this factor can be accounted in future studies. Combining all the terms, the photon transmission rate reads $\eta(z,\theta)=\eta_{\text{tran}}\eta_{\text{d}}(z)\eta_{\text{atm}}(\theta)\eta_{\text{rec}}\eta_{\text{det}}\eta_{\text{filt}}\eta_{\text{coup}}$, where the extra $\eta_{\text{coup}}$ is the single-mode optical fiber coupling efficiency enhanced by a high-bandwidth adaptive-optics system [35, 36].
@@
The photon source used on Micius operated with a rate of $5.9$ million entangled photon pairs per second and a fidelity of $0.907\pm 0.007$. The successfully detected photon pairs were still entangled with a fidelity of $0.869\pm 0.085$. Since noise from the environment is negligible when entangled photons are transmitted through the vacuum (before reaching the Earth’s atmosphere), a MEO satellite with the same source would distribute entanglement with the same fidelity. Due to the much larger serving area of MEO satellite, we assume the satellite is equipped with multiple such entangled photon sources, resulting in a total generation rate of $N_{\text{s}}=6\times 10^{7}$ with a fidelity similar to Micius’s, namely $F_{\text{s}}=0.87$.
IV IV. Hybrid satellite optical-fiber network
@@
To study the hybrid protocol for entanglement distribution combining both optical fibers and satellites in a large-scale quantum network, we first focus on the physical aspects of a representative case. We model a hybrid quantum network for the contiguous United States using MEO satellites and a ground optical-fiber network designed using the population distribution data from the 2020 national census [38]. The ground network has a total of 83 0478304783\,04783 047 nodes, each representing the latitude and longitude coordinates of the centroid of a census tract region (or the point within the region that is closest to the centroid). A census tract is a subdivision of a county that is used by the U.S. Census Bureau for organizing and analyzing census data, each including an average of 3 957 residents.
@@
Figure 3 shows the resulting network, where each node is assumed to be connected by two-way optical fibers (undirected links) to its three closest neighbors. In order to allow at least one path of optical fiber links connecting any two nodes (i.e., to form a single network connected component), an extra 24 links are added while minimizing the total fiber length [Fig. 3(a)], leading to a total of 162 860162860162\,860162 860 links. The number of neighbors connected to a node (i.e., the node degree) ranges from 3 to 9, with an average of $3.9$ [Fig. 3(b)].

### v2

@@ Abstract
Quantum networks hold promise for key distribution, private and distributed computing, and quantum sensing, among other applications. The scale of such networks for ground users is currently limited by one’s ability to distribute entanglement between distant locations. This can in principle be carried out by transmitting entangled photons through optical fibers or satellites. The former is limited by fiber-optic attenuation while the latter is limited by atmospheric extinction and diffraction. Here, we propose a hybrid network and protocol that outperform both ground- and satellite-based designs and lead to high-fidelity entanglement at a continental or even global scale.
@@
The ongoing development of quantum technologies calls for advances in quantum communication. The basis of quantum communication is entanglement, which, once established, can allow the transmission of quantum information over arbitrary distances [1]. Multiple small-scale quantum networks have been successfully implemented [2, 3, 4, 5, 6]. The prospect of developing large-scale quantum networks is motivated by applications ranging from quantum key distribution [7, 8] and distributed quantum computing [9] to quantum navigation and sensing [10, 11]. However, scalability remains a challenge due to difficulties associated with the generation of entanglement between distant parties, and thus new approaches are needed to lay the groundwork for the future development of a quantum internet [12, 13].
@@
The most direct approach for distributing entanglement in a quantum network is by using optical fibers to transmit entangled photons [14, 15]. Due to the exponential loss of photons with distance, establishing long-distance entanglement requires intermediate quantum repeaters [16, 17]. Such devices, which are currently under development, are able to generate entangled states, transmit/receive entangled photons, perform quantum operations, and/or store qubits. Within this approach, entanglement is first directly distributed over short distances. Long-distance entanglement is then established through entanglement swapping performed at intermediate quantum repeaters. However, existing quantum repeater technology currently limits entanglement distribution over long distances [18].
@@
An alternative approach is to use satellites as sources of entangled photon pairs that are sent to two distant detection stations on the ground [19]. Entanglement distribution is successful if both ground stations detect photons of the same pair [20]. The challenge with this approach is that photons are subject to diffraction loss, where the survival probability scales with the inverse square of the distance, and absorption and scattering in Earth’s atmosphere. In 2016, China launched Micius, the first quantum communication low Earth orbit (LEO) satellite [21], which was equipped with an entangled-photon-pair source and operated in an orbit 500 km from the ground. Entanglement distribution was established between two ground stations 1200 km apart from each other [20]. Moreover, this LEO satellite was successfully used to enable quantum key distribution between China and Europe that are 7600 km apart on Earth [22]. Subsequent work has also developed scheduling policies for entanglement distribution using groups of LEO satellites [23].
@@
While using LEO satellites to distribute entanglement results in low photon loss, certain challenges with this technology currently impact its scalability [21]. First, LEO satellites travel at high speeds relative to the ground, leaving only short time windows for interaction with the corresponding stations, and thus requiring high-precision acquisition, pointing, and tracking systems. Second, the short distances between LEO satellites and the ground limit the area that can be covered by a single satellite, making it impossible for a satellite to establish entanglement between stations that are far apart. This second shortcoming can be addressed by allowing the satellite to temporarily store qubits as a trusted node, but this could compromise security [24]. An alternative to mitigate both shortcomings is to use geostationary (GEO) satellites, positioned at 36 000 km above the equator in a fixed position relative to the ground [25]. A single GEO satellite can cover more than one-third of the Earth’s surface, but the loss caused by diffraction is $10^{4}$ times higher than for Micius, which limits the entanglement distribution rate.
@@
In this article, we investigate a hybrid quantum network consisting of a combination of optical fibers and medium Earth orbit (MEO) satellites. The goal is to benefit from the short-distance advantage of optical fibers and the long-distance advantage of satellites. Our focus on MEO satellites—chosen to be 10 000 km above ground—seeks to strike a compromise between low photon loss and high spatio-temporal coverage. Using the contiguous United States as a model system, we show that the hybrid protocol provides better large-scale quantum communication performance than both optical-fiber-based and satellite-based protocols.
@@
II II. Optical-fiber transmission
Because the number of photons successfully transmitted through an optical fiber decreases exponentially with distance, various quantum repeater designs have been proposed to compensate for this loss [17]. In particular, a promising trapped-ion quantum repeater has been recently demonstrated in an entanglement distribution experiment between points separated by 50 km of optical fiber [26]. A trapped-ion repeater includes at least two trapped ions and the ability to emit photons entangled with each ion. The trapped ions in the repeaters serve as quantum memories storing qubits. Entanglement swapping between the ions can then be implemented through deterministic Bell-state measurement (DBSM).
The ground portion of our hybrid quantum network will consist of a natural extension of this scheme, in which long-distance entanglement distribution is achieved using multiple quantum repeaters [Fig. 1(a)]. In this extended scheme, a photon repeater is placed between every pair of trapped-ion repeaters (including the end nodes) [26]. The photon repeaters can detect photons and perform quantum entanglement swapping using probabilistic Bell-state analysis. Alice and Bob (the end nodes) are assumed to have one trapped ion and, as trapped-ion repeaters, the ability to emit ion-entangled photons. Thus, the system consists of two end nodes, $n_{\text{p}}$ photon repeaters, and $n_{\text{i}}=n_{\text{p}}-1$ trapped-ion repeaters for $n_{\text{p}}\geq 1$ ($n_{\text{i}}=0$ otherwise). In this system, ion-photon entangled states $\rho_{\text{ip}}$ are converted into ion-ion entangled states $\rho_{\text{ii}}$ and then to an end-to-end entangled state $\rho_{\text{AB}}$ shared by Alice and Bob.
Figure 2: Entanglement distribution rate and fidelity through optical fibers. (a) Distribution rate without repeaters (black) and for an optimal number of evenly placed photon (and thus intermediate trapped-ion) repeaters (blue) as a function of the optical fiber length $L$. The optimal number of photon repeaters is marked on the top axis. Inset: magnification of both curves for small $L$. (b) Resulting end-to-end fidelity as the number of evenly placed repeaters is varied. (c) Corresponding entanglement distribution rate for two fixed values of $L$ as the number of repeaters is varied.
@@
where $L$ is the fiber length between the end nodes, and the other symbols are defined as follows. The terms in parentheses account for the time $\tau$ needed to emit an entangled photon and the time to transmit the photon to an adjacent photon repeater, where $c=2\times 10^{8}$ m/s is the speed of light in the optical fiber; note that a factor of 1/2 is canceled because the trapped-ion nodes wait to hear back from photon repeaters whether the photon was detected or not before attempting to send a new photon. The term $\frac{1}{2}P^{2}\eta^{2}$ is the probability that both photons concurrently directed to a photon repeater are successfully received, detected, and projected into a Bell state; here, $\eta(L)=10^{-\gamma L/2n_{\text{p}}}$ is the fiber transmission rate for an attenuation rate $\gamma$, $P$ is the photon-detector efficiency, and 1/2 is the probability of a successful photon Bell-state measurement. The waiting time to establish a successful entanglement between one pair of adjacent trapped-ion nodes follows a geometric distribution; thus, the additional factor $(3/2)^{\nu}$ is a good approximation for the waiting time before all pairs in the path are ready to be entangled, where $\nu=\log_{2}{n_{\text{p}}}$. The time for local quantum operations associated with entanglement swapping is assumed to be negligible, with entanglement swapping having a success rate of 1/2 for photon repeaters (as mentioned) and 1 for trapped-ion repeaters. The average entanglement distribution rate is then given by $R_{\text{f}}=1/T_{\text{f}}$. As a reference, we note that the average time required to establish an entanglement between Alice to Bob without any intermediate repeater is
@@
where $\eta^{\prime}(L)=10^{-\gamma L}$, and the average entanglement distribution rate is thus given by $R_{\text{f}}^{\prime}=1/T_{\text{f}}^{\prime}$.
[…]
A competing factor to be accounted for is the end-to-end fidelity as the number of repeaters increases. We follow the model discussed in [28] to calculate the fidelity as a function of $n_{\text{p}}$. We assume that the initially generated ion-photon entanglement is a mixed state of the form $\rho_{\text{ip}}=F_{0}|\Psi^{-}\rangle\langle\Psi^{-}|+(1-F_{0})|\Psi^{+}\rangle\langle\Psi^{+}|$, where $|\Psi^{\pm}\rangle$ are Bell states and $F_{0}$ is the pairwise fidelity [Fig. 1(a)]. A successful Bell-state measurement on a photon pair, which occurs with a probability $1/2$, leads to a new mixed entangled state $\rho_{\text{ii}}=F_{\text{ii}}|\Psi^{-}\rangle\langle\Psi^{-}|+(1-F_{\text{ii}})|\Psi^{+}\rangle\langle\Psi^{+}|$ shared by two neighboring trapped-ion nodes. The ion-ion entanglement fidelity is $F_{\text{ii}}=[1+V(1-2F_{0})^{2}]/2$, where $V=2F_{\text{swap}}^{\text{p}}-1$ is the photon interference-visibility. Each intermediate trapped-ion repeater node can then perform a DBSM on the two trapped-ion qubits they maintain. Given $\rho_{\text{i}_{1}\text{i}_{2}}$ and $\rho_{\text{i}_{3}\text{i}_{4}}$, an ideal DBSM will give $\rho_{\text{i}_{1}\text{i}_{4}}$ by projecting trapped-ions 2 and 3 onto a Bell state and tracing them out. If we account for the depolarizing noise caused by the DBSM, the resulting quantum state will be $\tilde{\rho}_{\text{i}_{1}\text{i}_{4}}(\rho_{\text{i}_{1}\text{i}_{4}},F^{\text{i}}_{\text{swap}})$, where $F_{\text{swap}}^{\text{i}}$ is the trapped-ion entanglement-swapping fidelity and
@@
Here, $S_{j}\in[I\otimes\sigma_{x},I\otimes\sigma_{y},I\otimes\sigma_{z}]$, where $\sigma_{j}$ are the Pauli matrices and $I$ is the identity matrix. The fidelity of the final end-to-end entangled state is obtained after completing all the entanglement-swapping operations on the intermediate trapped-ion nodes. Through the entire process, the fidelity of the established entanglement can decay due to the decoherence of qubits stored in the quantum memories. For an arbitrary initial state $\rho_{0}$, this can be modeled by considering dephasing noise of the form
@@
where $q(t)=(1-e^{-t^{2}/\tau_{q}^{2}})/2$ and $\tau_{q}$ is the decoherence time. A recent study showed that the decoherence time of trapped-ion entanglement can be longer than 10 s [29], which is much longer than the relevant time needed to establish one end-to-end entanglement discussed in this study. Therefore, we assume the impact of decoherence to be negligible in our calculations, leading to no fidelity loss. An analysis of the impact of non-negligible decoherence is included in Appendix D.
[…]
Figure 3: Model optical-fiber network of the contiguous U.S. (a) Network structure, where each node represents a census tract and each link represents an optical fiber. Links that require intermediate repeaters, namely those longer than 61.7 km, are marked green. The orange path illustrates an entanglement distribution route between end nodes Alice and Bob. (b-d) Histograms of node degrees (b), link lengths (c), and optical-fiber path lengths between 10 00010\,000 randomly sampled end-node pairs (d) in the network.
@@
We consider an MEO satellite orbiting 10 000 km above ground. The satellite is equipped with an entangled-photon source that can generate $N_{\text{s}}$ entangled photon pairs per second. The photon pairs are split into two collimated beams and are then sent toward two ground stations as shown in Fig. 1(b). An entanglement is successfully established between two ground stations if both stations detect photons from the same entangled pair. The ground stations will then store the quantum states of the received photons in single-photon quantum memories [which also use trapped ions, but are different from the trapped-ion memories that operate by emitting photons in Fig. 1(a)] [31]. Free-space diffraction and atmospheric extinction (absorption and scattering) are accounted for as the main factors causing losses during transmission [32]. Other losses come from limitations in the transmitting and receiving telescopes.
Figure 4: Entanglement distribution efficiency through the exclusive use of optical fibers or MEO satellites for 10 000 randomly sampled Alice-and-Bob pairs in the network shown in Fig. 3(a). (a) Distribution rate through optical fibers (blue) and a satellite (orange). For a fair comparison, the fidelity for the satellite and target fidelity for the ground network are both assumed to be $0.87$. (b) Fidelity of the end-to-end entanglement for the time needed to establish one entanglement via satellite, color-coded as in (a). The points $D^{\star}_{\text{R}}$ and $D^{\star}_{\text{F}}$ mark the distances where the relative efficiency inverts. Repeaters are placed to achieve optimal distribution rate in (a) and optimal fidelity in (b). (c) Scattered plot (purple) and histogram (grey) of the distribution rate in (a) and fidelity in (b).
@@
$w_{\text{d}}(z)=w_{0}\sqrt{(1+z^{2}/z_{\text{R}}^{2})},$ (5)
where $z_{\text{R}}=\pi w_{0}^{2}\lambda^{-1}$ is the Rayleigh range, $\lambda$ is the wavelength, and $w_{0}$ is the initial beam width. The slant range $z$ can be easily calculated given the longitudes and latitudes of the ground station and the subsatellite point (the projection point of the satellite on the Earth’s surface) by converting them to Cartesian coordinates. Given a telescope receiver of a radius $a_{\text{r}}$, the fraction of the initial beam that can be detected is
@@
for the far field $z\gg z_{\text{R}}$ relevant for MEO satellites. Absorption and Rayleigh/Mie scattering occur when the beam travels through the atmosphere (mainly the $30$ km layer above the ground). The fraction of the beam surviving absorption and scattering is
@@
where $\eta(z,\theta)$ is the photon transmission rate between the satellite and one ground station. This rate accounts for the efficiency $\eta_{\text{tran}}$ of the transmitting telescope and the optics efficiency $\eta_{\text{rec}}$, detection efficiency $\eta_{\text{dec}}$, and spectral-filter efficiency $\eta_{\text{filt}}$ of the receiving telescope at the ground station [33, 34]. The photons also need to be converted to a wavelength and spatiotemporal mode compatible with the quantum memories at the ground stations. Since the conversion efficiency can vary depending on the type of memory used, we assume a perfect conversion for simplicity; this factor can be accounted for in future studies. Combining all the terms, the photon transmission rate reads $\eta(z,\theta)=\eta_{\text{tran}}\eta_{\text{d}}(z)\eta_{\text{atm}}(\theta)\eta_{\text{rec}}\eta_{\text{det}}\eta_{\text{filt}}\eta_{\text{coup}}$, where the extra $\eta_{\text{coup}}$ is the single-mode optical-fiber coupling efficiency enhanced by a high-bandwidth adaptive-optics system [35, 36].
@@
The photon source used on Micius operated with a rate of $5.9$ million entangled photon pairs per second and a fidelity of $0.907\pm 0.007$. The successfully detected photon pairs were still entangled with a fidelity of $0.869\pm 0.085$. Since noise from the environment is negligible when entangled photons are transmitted through the vacuum (before reaching the Earth’s atmosphere), an MEO satellite with the same source would distribute entanglement with the same fidelity. Due to the much larger serving area of MEO satellite, we assume the satellite is equipped with multiple such entangled photon sources, resulting in a total generation rate of $N_{\text{s}}=6\times 10^{7}$ with a fidelity similar to Micius, namely $F_{\text{s}}=0.87$.
IV IV. Hybrid satellite and optical-fiber network
@@
To study the hybrid protocol for entanglement distribution combining both optical fibers and satellites in a large-scale quantum network, we first focus on the physical aspects of a representative case. We model a hybrid quantum network for the contiguous United States using MEO satellites and a ground optical-fiber network designed using the population-distribution data from the 2020 national census [37]. The ground network has a total of 83 04783\,047 nodes, each representing the latitude and longitude coordinates of the centroid of a census tract region (or the point within the region that is closest to the centroid). A census tract is a subdivision of a county that is used by the U.S. Census Bureau for organizing and analyzing census data; each includes an average of 3 957 residents.
@@
Figure 3 shows the resulting network, in which each node is assumed to be connected by two-way optical fibers (undirected links) to its three closest neighbors. To allow at least one path of optical-fiber links connecting any two nodes (i.e., to form a single network connected component), an extra 24 links are added while minimizing the total fiber length [Fig. 3(a)], leading to a total of 162 860162\,860 links. The number of neighbors connected to a node (i.e., the node degree) ranges from 3 to 9, with an average of $3.9$ [Fig. 3(b)].